OpenStudy (anonymous):

the regular price of a child's entry ticket to a water park is $8 less than for an adult's. the park offers half off all entry tickets during the off- peak season. The Sandlers paid a total of 194 for 1 adult ticket and 4 child tickets to the water park during the off- peak season the following equation represents this situation, where x represents the regular price of an adult ticket 194= 1/2x + 2(x-8 What is regular price of a child's ticket? $68 $76 $84 $89

OpenStudy (anonymous):

194 = 1/2x + 2(x - 8) -- distribute through the parenthesis 194 = 1/2x + 2x - 16 194 + 16 = 1/2x + 2x 210 = 1/2x + 4/2x 210 = 5/2x 210 / (5/2) = x 210 * 2/5 = x 420/5 = x 84 = x any questions ?